OpenStudy (anonymous):

ok guys i really need help on literal equations! y= 3/4x+9 (solve for x)

OpenStudy (anonymous):

4xy+9y=3 4xy=3-9y x=(3-9y)/4y

OpenStudy (anonymous):

x=(3-9y)/4y